A nut mix contains 10% of peanuts and 90% of cashews. It costs 5% more than pure
peanuts. Are cashews more expensive than peanuts? By how much?

Found 2 solutions by Jk22, onlinepsa:
Answer by Jk22(389)

Since with cashew it costs more than pure peanuts, cashews have to be more expensive.


let p the price of peanuts, and c of cashews, we have

0.1p*+0.9c=1*p+0.05*p

the price of 10% peanuts and 90% cashews is the price if it were peanuts plus 5% of it.

Express c as function of p :

0.95p = 0.9c => c = .95/.9p = 1.05555p,


the price of cashews is the one of peanuts +5.555% = +50/9 %

Answer by onlinepsa(22)

Let peanuts cost 'x' for 100 gm.
A nut mix of 100 gm contains 10 gm of peanuts and 90 gm of cashews.
Cost of nut mix of 100 gm= x + 5*x/100

Also, Cost of nut mix= Cost of 10 gm of peanut + cost of 90 gm of cashews
=>Cost of nut mix= (x/10) + cost of 90 gm of cashews
=>x + x/20 = (x/10) + cost of 90 gm of cashews
=>cost of 90 gm of cashews= x + x/20 - (x/10)
=>cost of 90 gm of cashews= 19x/20
=>cost of 10 gm of cashews= 19x/9*20 = 19x/180
=>cost of 100 gm of cashews= 190x/180= 19x/18= (1+1/18)x

Cashews are 5.56% more costly than peanuts.